In the context of dividing a bag of 18 oranges into sixths, explain what is the "whole" which is shared equally into six parts? Now answer the question: in which case may s " whole" not be one?


1
Expert's answer
2021-06-07T17:09:49-0400

"whole"=18/6=3


" whole" is not one:

dividing a bag of 18 oranges into 2 parts:

"whole"=18/2=9


dividing a bag of 18 oranges into 3 parts:

"whole"=18/3=6


dividing a bag of 18 oranges into 2 parts:

"whole"=18/9=2